Joshua visits the computer shop every 6 days while Jacob visits the same place every 8 days. If they both visit the computer shop on July 1, what is the earliest date when they will both visit the computer shop again?
Answer by nerdybill(7384)

Joshua visits the computer shop every 6 days while Jacob visits the same place every 8 days. If they both visit the computer shop on July 1, what is the earliest date when they will both visit the computer shop again?
.
Find the LCM of 6 and 8:
factor each using prime number:
6: 2 3
8: 2 2 2
.
LCM = 2*2*2*3 = 24
.
July 1 + 24 days = July 25